HUBBLE SPACE TELESCOPE DAILY REPORT #1375 PERIOD COVERED: 0000Z (UTC) 04/21/95 - 0000Z (UTC) 04/24/95 Daily Status Report as of 114/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Set of WF/PC-2 5463 (Buried QSOs in IRAS-Selected Ultra luminous AGNs) The WF/PC-2 was used to make observations of the luminous IRAS active galactic nuclei of galaxies IRAS0910+4 and IRAS2046+1 to try and resolve the polarized scattering regions in their cores. The first set of observations were completed as planned, and no anomalies were reported. However, during the second set of observations the guide star acquisition failed to coarse track at 113/1644Z due to a scan step limit being exceeded. The following reacquisitions also failed to coarse track. The total time spent in coarse track was approximately five hours. The problem appeared to be a bad guide star. 1.2 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.3 Completed WF/PC-2 5401 (Luminosities and Morphologies of a Complete 3CR Sample of Radio Galaxies) The WF/PC-2 was used to obtain images of radio galaxy 3C175.1, which has a Z near 1.0, and also is a quasar. These observations will be used to determine and compare luminosities, and morphologies in the optical and infrared. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Three Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 5393 (The Nature of the Host Galaxies of High-Redshift, Radio-Loud Quasars) The WF/PC-2 was used to make observations of galaxy MRC0549-21 which contains a high red-shifted, radio-loud quasar to study its morphology. This data will be compared with low-Z QSOs and other radio galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 5132 (A Photometric Search for Planets of Nearby Stars) The WF/PC-2 was used to make photometric observations of nearby star BD+4D2561H to search for possible planets. The exposure time was calculated so that area beyond 1.2 arcseconds from the central star is not saturated. The observations were completed as planned, and no problems were reported. 1.7 Completed Two Sets of FOS/BL 6146 (FOS Location of Spectra and Aperture - Part 2) The Faint Object Spectrograph (Blue Side) was used to conduct a test to determine the repeatability of positioning the filter-grating wheel. All observations were internal. A wave length calibration image, Taled calibration and a Dark image were also taken. All observations were executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 5106 (Planetary Nebular Structure (WC03): Cycle 4) The WF/PC-2 was used to make images of the planetary nebula NGC-7293 to study long lived condensations that appear in this object. The observations were executed as planned, and no problems were reported. 1.9 Completed WF/PC-2 5629 (WF/PC-2 Cycle 4 Photometric Calibration Monitor 1C: 4-Chip STD in F170W) The WF/PC-2 was used to make photometric calibration observations of the ultraviolet standard star GRW+70D582. The guide star acquisition at 112/2005Z failed during ZOE (HSTAR 5187). The search radius limit was exceeded, and the following update roll delay indicated a V3 error of 86 arcseconds. There was no time to schedule an ARU/PRT corrective maneuver, and this effected the next set of observations (See Note 1.10). 1.10 Completed FOS/BL 5459 (Ultraviolet Interstellar Polarization: Cycle 4 High) The Faint Object Spectrograph (Blue Side) was used to make ultraviolet spectrographic observations of the bright star HD207538 to study interstellar polarization along the line of sight to the star. This is to study interstellar grains that are responsible for the extinction and polarization of the star light. The previous observations encountered a guide star acquisition failure (See Note 1.9), and there was no time to unlink a corrective ARU/PRT maneuver. Therefore, the acquisition at 112/2101Z failed (HSTAR 5187). The ARU was successfully uplinked to the spacecraft at 112/2246Z and the subsequent acquisition at 112/2350Z was successful. 1.11 Completed WF/PC-2 5371 (HST Medium Deep Survey: Cycle 4 Medium Low Galactic Latitude) The WF/PC-2 was used in parallel with the prime science observations to make low galactic latitude deep multicolor images of objects such as galaxies, and stars in our galaxy. The observations were completed as planned, and no problems were reported. 1.12 Completed WF/PC-2 5450 (Host Galaxies and Cluster Environments of Quasars) The WF/PC-2 was used to make observations of Quasar Q1222+1010 to study the properties of its host galaxy and surrounding environment. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of F294 to study the interactions between the central engine, the stars, and the interstellar matter in the AGN. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 16 Successful : 15 There was a guide star acquisition failure at 112/2005Z (During a ZOE). Scheduled Reacquisitions : 19 Successful : 19 2.2 FHST Updates: Scheduled : 48 Successful : 48 2.3 Significant HSTARs: None 2.4 Operations Notes: Last week, the bit slips while dumping ESTR #2 continued, but at a greatly reduced rate. Also, there were data inversions while operating from STGT, but only on three occasions on Sunday. FSW 9.5 was installed completely by 113/1228Z. There were no reported problems. The CoP status is being monitored by TALOS. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-